### v1

# Fine-grained Image Retrieval via Dual-Vision Adaptation

Abstract.
Fine-Grained Image Retrieval (FGIR) faces challenges in learning discriminative visual representations to retrieve images with similar fine-grained features. Current leading FGIR solutions typically follow two regimes: enforce pairwise similarity constraints in the semantic embedding space, or incorporate a localization sub-network to fine-tune the entire model. However, such two regimes tend to overfit the training data while forgetting the knowledge gained from large-scale pre-training, thus reducing their generalization ability. In this paper, we propose a Dual-Vision Adaptation (DVA) approach for FGIR, which guides the frozen pre-trained model to perform FGIR through collaborative sample and feature adaptation. Specifically, we design Object-Perceptual Adaptation, which modifies input samples to help the pre-trained model perceive critical objects and elements within objects that are helpful for category prediction. Meanwhile, we propose In-Context Adaptation, which introduces a small set of parameters for feature adaptation without modifying the pre-trained parameters. This makes the FGIR task using these adjusted features closer to the task solved during the pre-training. Additionally, to balance retrieval efficiency and performance, we propose Discrimination Perception Transfer to transfer the discriminative knowledge in the object-perceptual adaptation to the image encoder using the knowledge distillation mechanism. Extensive experiments show that DVA has fewer learnable parameters and performs well on three in-distribution and three out-of-distribution fine-grained datasets.
Fine-grained image retrieval, dual vision adaptation, knowledge distillation
copyright: none
1. Introduction
Unlike general image retrieval, fine-grained image retrieval (FGIR) attempts to retrieve images belonging to the same subcategory as the query image from a database within a broader meta category (i.e., birds, cars) (Wei et al., 2017). It has been extensively applied across various domains, including intelligent transportation (Ramdurai, 2025) and biodiversity monitoring (Vendrow et al., 2024). However, FGIR presents significant challenges due to two inherent difficulties: i) subtle visual distinctions between different categories, and ii) significant appearance variations within the same category. This task usually requires models to simultaneously localize discriminative regions and identify minute visual differences, creating a dual optimization dilemma. Therefore, the key to FGIR is learning discriminative and generalizable embeddings to identify visually similar objects accurately.
Conventional FGIR methods typically adopt two primary strategies: encoding-based (Movshovitz-Attias et al., 2017; Teh et al., 2020; Ermolov et al., 2022; Lim et al., 2022) and localization-based (Zheng et al., 2018; Wang et al., 2023b; Wei et al., 2017; Wang et al., 2022). Encoding-based methods primarily learn image-level features but struggle to suppress background and irrelevant information interference. In contrast, localization-based methods enhance feature extraction within the encoder to capture subtle differences among subcategories, often by focusing on distinctive object regions. Although both strategies capture discriminative embedding learning for fine-grained retrieval, their dependence on full-parameter fine-tuning introduces prohibitive computational costs and limits cross-task generalization, as shown in Fig. 1(a). Simultaneously, fine-tuning the entire model destroys the knowledge in the large-scale pre-trained parameters and may lead to suboptimal convergence, which ultimately limits retrieval performance (Wang et al., 2023b). This raises the question: can discriminative representations be learned for FGIR tasks without fine-tuning the model’s pre-trained parameters?
Figure 1. (a) previous fine-tuning methods. (b) our dual-vision adaptation method. Our approach designs the collaborative sample and feature adaptation to exploit category-specific differences. This dual strategy enables model to sustain broad representation capabilities from pre-training data while dynamically adjusting its adaptability to fine-grained data.
Recent advances in parameter-efficient fine-tuning (PEFT) techniques have demonstrated the feasibility of learning representations while introducing a limited number of new parameters to adapt a frozen pre-trained model to downstream tasks. Its core idea is to redesign downstream tasks to align closely with those addressed during pre-training, ensuring that large-scale pre-trained knowledge is effectively retained. Owing to PEFT, large-scale pre-trained visual-language models (e.g., CLIP (Radford et al., 2021)) have achieved impressive results in various visual tasks. However, existing PEFT methods (Zhou et al., 2022a, b) are designed for multimodal models to capture high-level category semantics rather than subtle subcategory differences, so directly applying existing PEFT techniques to FGIR without task-specific adaptation tends to be suboptimal, due to their limited focus on capturing subtle intra-class variations. Thus, developing an efficient fine-tuning strategy tailored for fine-grained visual models is essential to avoid suboptimal convergence when fine-tuning the entire FGIR model.
To this end, we design a novel adaptation framework for FGIR, which efficiently unleashes the power of pre-trained model, as shown in Fig. 1(b). Technically, we propose Dual-Vision Adaptation (DVA) to resolve the conflict between preserving pre-trained knowledge and acquiring fine-grained discriminative power. By maintaining the frozen backbone, DVA establishes category-sensitive perception through sample-feature co-adaptation, inherently avoiding the optimization dilemma of conventional full fine-tuning. Specifically, the Object-Perceptual Adaption (OPA) is a proposed sample-adapted strategy for the visual foundation model, encouraging the pre-trained model to capture critical object regions and locate subclass-specific differences. Meanwhile, as a feature adaptation process, the proposed In-Context Adaption (ICA) aims to focus on fine-grained features that contribute to category prediction through lightweight learnable parameters. In this way, ICA can adapt the feature content in a direction that is conducive to category recognition, which makes the FGIR task with such adjusted features close to the task solved during the original pre-training. Nevertheless, a non-negligible problem is that the integration between the visual foundation model and the frozen pre-trained backbone introduces prohibitive computational overhead during inference. To address this critical efficiency bottleneck while retaining fine-grained discriminative power, we devise a Discriminative Perceptual Transfer (DPT) module that transfers the discriminative knowledge from OPA to the backbone via a distillation mechanism.
In summary, the primary contributions of this work are as follows:
  • We propose a dual vision adaptation framework to enable the frozen pre-trained model to capture subtle sub-category differences without forgetting the pre-trained knowledge, balancing accuracy and efficiency.
  • We propose a dynamic sample and feature co-adaptation strategy and a distillation-based sample perception transfer module to bridge the gap between high-level semantics and fine-grained distinctions, which are realized through the collaborative modification of samples to features.
  • With only 3.5% of the tunable parameters compared to full fine-tuning, DVA achieves competitive performance on three in-distribution and three out-of-distribution fine-grained datasets.
2. Related Works
2.1. Fine-Grained Image Retrieval
Existing methods for fine-grained image retrieval (FGIR) can be categorized as encoding-based or localization-based. The encoding-based methods (Movshovitz-Attias et al., 2017; Teh et al., 2020; Lim et al., 2022; Ermolov et al., 2022; Jiang et al., 2024c, a) aim to learn an embedding space in which samples of a similar subcategory are attracted and samples of different subcategories are repelled. The methods can be decomposed into roughly two components: the image encoder maps images into an embedding space, and the metric method ensures that samples from the same subcategories are grouped closely, while samples from different subcategories are separated. While these studies have achieved significant achievements, they primarily concentrate on optimizing image-level features that include numerous noisy and non-discriminatory details. Therefore, the localization-based methods (Han et al., 2018; Tang et al., 2020; Wang et al., 2023b; Wei et al., 2017; Moskvyak et al., 2021; Wang et al., 2022) are proposed, which focus on training a subnetwork for locating discriminative regions or devising an effective strategy for extracting attractive object features to facilitate the retrieval task. Unlike these approaches, our method considers the specific characteristics of the FGIR tasks, offering guidance for designing high-performance retrieval models.
2.2. Parameter-Efficient Fine Tuning
Parameter-efficient fine-tuning (Hu et al., 2022) in NLP reformulates downstream tasks as language modeling problems, allowing pre-trained language models to adapt more efficiently to new tasks. As a result, these techniques are now widely used in various NLP applications, including language understanding and generation. Recently, parameter-efficient fine-tuning has also been applied to multimodal computer vision (Zhou et al., 2022b, a; Jiang et al., 2024b; Cao et al., 2023, 2024b, 2024a). However, existing methods mainly extend the capabilities of language-based models, making them unsuitable for direct use with pre-trained vision models. Additionally, some recent approaches (Chen et al., 2022a; Gao et al., 2024) introduce a small number of learnable parameters to guide pre-trained models in general vision tasks. However, their fine-tuning strategies focus on capturing category-level semantics rather than the fine-grained visual differences needed to distinguish similar objects. To address this limitation, we propose a parameter-efficient DVA that incorporates feature and sample adaptation, enabling the frozen pre-trained vision model to perform FGIR tasks effectively.
2.3. Knowledge Distillation
Knowledge Distillation (KD) is dedicated to compressing the informative knowledge from a large and computationally expensive model (i.e, teacher model) to a small and computationally efficient model (i.e, student model) (Hinton et al., 2015; Fang et al., 2024). Given the strong knowledge transfer ability, it is widely applied to natural language processing, computer vision and other fields (Wang and Yoon, 2021; Hu et al., 2023). Most classification-based KD approaches explore enhancing the student model by mimicking the teacher model’s predictions or output distribution. Additionally, researchers have delved into exploring KD for image retrieval by leveraging distances between samples, such as learning to rank (Chen et al., 2018) and regression on quantities involving one or more pairs, like distances (Wang et al., 2023a) or angles (Park et al., 2019). In this paper, we transfer discriminative knowledge that aids fine-grained image retrieval via proxy features.
3. Methods
Figure 2. The overall framework of our proposed DVA. It consists of three crucial components, $i.e$, Object-Perceptual Adaption (OPA), In-Context Adaption (ICA) and Discriminative Perceptual Transfer (DPT). OPA enhances the image encoder’s ability to capture discriminative object regions and ICA dynamically adapts fine-grained features while suppressing irrelevant background retained in the frozen representations. Finally, DPT inject a discriminative awareness into the image encoder via distillation, enabling auxiliary-free inference while preserving pre-trained knowledge.
3.1. Image Encoder
For the input image $\textbf{I}\in\mathbb{R}^{3\times H\times W}$, the image encoder will first split it into $N=N_{h}\times N_{w}$ non-overlapping patches of size $P\times P$. Here, $N_{h}=\frac{H}{P}$ and $N_{w}=\frac{W}{P}$. Subsequently, these patches are transformed into embedding tokens $\textbf{E}=[\textbf{E}^{1},\textbf{E}^{2},\dots,\textbf{E}^{N}]\in\mathbb{R}^{N\times D}$ using learnable linear projection, where $D$ denotes the dimension of each token. Finally, the embedding tokens E are concatenated with a class token $\textbf{E}^{\mathrm{{cls}}}\in\mathbb{R}^{D}$, and the position information of the tokens is retained through combined a learnable position embedding $\textbf{E}_{\mathrm{pos}}$ to form the initial input token sequence as $\textbf{E}_{0}=[\textbf{E}^{\mathrm{cls}},\textbf{E}^{1},\textbf{E}^{2},\dots,\textbf{E}^{N}]+\textbf{E}_{\mathrm{pos}}$. The image encoder comprises $L$ layers, each incorporating multi-head self-attention (MHSA) and a multi-layer perception (MLP) block (Dosovitskiy et al., 2020). Therefore, considering the input $\textbf{E}_{i-1}$ of the $i$-th layer, the output $\textbf{E}_{i}$ can be obtained as follows:
(1) $\mathbf{E}_{i}^{\prime}$ $=\textrm{MHSA}(\textrm{LN}(\mathbf{E}_{i-1}))+\mathbf{E}_{i-1},$
(2) $\mathbf{E}_{i}$ $=\textrm{MLP}(\textrm{LN}(\mathbf{E}_{i}^{\prime}))+\mathbf{E}_{i}^{\prime},$
where $i=\{1,2,\dots,L\}$, and $\textrm{LN}(\cdot)$ denotes the layer normalization operation. More specifically, the MHSA module projects $\mathbf{E}_{i-1}$ into queries $\mathbf{Q}_{i-1}$, keys $\mathbf{K}_{i-1}$, and values $\mathbf{V}_{i-1}$, each $\in\mathbb{R}^{(N+1)\times D}$, with a attention projector.
(3) $\mathbf{Q}_{i-1},\mathbf{K}_{i-1},\mathbf{V}_{i-1}=\textrm{AttentionProjector}(\textrm{LN}(\mathbf{E}_{i-1})).$
These projections are then split into $H$ heads to achieve parallel attention, followed by an output projector to obtain the output $\mathbf{E}^{mhsa}_{i-1}$ of MHSA:
(4) $\mathbf{E}^{mhsa}_{i-1}=\textrm{OutputProjector}(\mathrm{softmax}\Bigl{(}\frac{\mathbf{Q}_{i-1}\mathbf{K}_{i-1}^{\top}}{\sqrt{D_{h}}}\Bigr{)}\mathbf{V}_{i-1}),$
where $D_{h}=\frac{D}{H}$.
Finally, we serve the class token $\textbf{E}^{\mathrm{cls}}$ from $\textbf{E}_{L}$ at the $L$-th layer as the retrieval embedding $\textbf{E}^{R}$.
3.2. Dual-Vision Adaptation
3.2.1. Object-Perceptual Adaption
Subtle yet discriminative discrepancies are widely recognized to be significant for fine-grained understanding (Fang et al., 2024; Wang et al., 2023b; Chen et al., 2022b). However, the vanilla ViT model is originally designed to identify different species (e.g., cats, dogs, and birds), rather than to exploit subtle differences between subcategories within a species. Therefore, without task-specific adaptation, the vanilla ViT has limited focus on capturing subtle intra-class variations. To alleviate this situation, we design an Object-Perceptual Adaption (OPA) that only modifies the visual input of the image encoder (ViT) to assist the model in perceiving discriminative differences in fine-grained images. Specifically, OPA consists of object perception and discriminative perception. The former better perceives objects through the background separated by the visual foundation model, and the latter perceives category differences through the critical foreground separated by the visual foundation model.
[…]
Discriminative Perception. For fine-grained understanding, there have been many works (Wang et al., 2023b; Jiang et al., 2024c) demonstrating that the background can perturb the model’s perception of discriminative regions. Therefore, in contrast to object perception, discriminative perception focuses on objects in the approach image, thereby helping to explore discriminative features through critical regions. We reuse the object perception process, the difference is that we remove the background area and keep the core object foreground, obtaining the discriminative image $\mathbf{I}_{\mathrm{d}}$. To ensure objects are fully contained in the image and free from deformation, we follow DVF (Jiang et al., 2024c) to filter out low-confidence detection results, and finally apply post-processing to preserve the original shapes of the object.
3.2.2. In-Context Adaptation
Fine-Tuning Preferences Analysis of ViT: Conventional methods (Ermolov et al., 2022; Fang et al., 2024) usually require fine-tuning the entire network, but this approach ignores the large-scale pre-trained knowledge present in the pre-trained parameters. We hypothesize that fine-grained image retrieval on downstream datasets can be learned and understood by specific modules within the network. To verify this hypothesis, we conducted an in-depth analysis of the fine-tune preferences of ViT.
We first divided the trainable parameters of transformer layers into three parts: the attention projector, output projector and MLP. Subsequently, we separately fine-tune these three parts and ViT as a whole, and use the fine-tuned models to test the results on the CUB-200-2011 (Branson et al., 2014) dataset. As shown in Table 1, we observed that only by fine-tuning part modules in ViT, we can achieve performance similar to that of fine-tuning the entire model. Meanwhile, fine-tuning the attention projector yields the best results. This is intuitive since the self-attention mechanism is central to ViT, and the attention projector directly determines output. This experiment further validates our hypothesis that the image retrieval task can be learned by a specific network module, namely attention projector.
Based on this insight, we attempt to tune only the attention projector of the ViT. However, this approach still destroys the knowledge in the pre-trained parameters and still requires learning a large number of parameters (22.5M). To address this issue, we keep the attention projector frozen and propose a lighter-weight learnable in-context adaptation module in parallel with the attention projector to learn the specific knowledge required for fine-grained image retrieval. Specifically, the in-context adaptation module is designed to be a bottleneck structure for limiting the number of parameters purpose, which includes a down-projection layer with parameters $\mathbf{W}_{down}\in\mathbb{R}^{D\times d}$, an up-projection layer with parameters $\mathbf{W}_{up}\in\mathbb{R}^{d\times D}$, in this work, we set $d=16$. For a specific input feature $\mathbf{x}$, the in-context adaptation module produces the in-context features, $\mathbf{x}_{ic}$, formally via:
(5) $\mathbf{x}_{ic}=\textrm{LN}(\mathbf{x})\cdot\mathbf{W}_{down}\cdot\mathbf{W}_{up}.$
Then, we can replace the original attention map output with a feature that contains both pre-trained knowledge and dataset-specific knowledge through residual connections:
(6) $\mathbf{\hat{Q}},\mathbf{\hat{K}},\mathbf{\hat{V}}=\textrm{AttentionProjector}(\textrm{LN}(\mathbf{E}))+\textrm{IC}(\mathbf{E}),$
where $\textrm{IC}(\cdot)$ represents the multiple in-context adaptation modules of $\mathbf{E}$ corresponding to the attention projectors that generate $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$. In this paper, we only assign in-context adaptation modules to the attention projectors that generate $\mathbf{Q}$ and $\mathbf{K}$. For more experimental analysis, please see Section 4.4.3.
Table 1. Fine-Tuning Preferences Analysis of ViT on the CUB-200-2011 dataset.
Methods R@1 R@2 R@4 R@8 Prams (M)
Fine-tuning 82.7 88.7 92.5 95.2 85.6
Attention Projector 82.7 89.0 92.8 95.2 21.5
Output Projector 82.6 88.8 92.6 95.0 7.3
MLP 82.0 88.9 92.9 95.2 56.8
3.3. Discriminative Perceptual Transfer
With the cooperation of discriminative perception, the fine-grained understanding ability of the ViT model is enhanced. However, the extra visual foundation model is time-consuming and memory-demanding for retrieval evaluation. In the classification field (Fang et al., 2024), network distillation has been shown to be one of the solutions to this problem. Inspired by network distillation, we propose a discriminative perceptual transfer to extend the knowledge distillation theory to retrieval tasks. Specifically, we learn discriminative category proxies as carriers of discriminative knowledge and transfer their discriminative knowledge using the same learning strategy.
Category Proxy Learning. We encode both the discrimination and distribution of labelled instances via proxy-guided learning (Kim et al., 2020). Then, for the retrieval embedding $\textbf{E}^{R}$ , we compute the proxy learning loss as follow:
(7) $\mathcal{L}_{proxy}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right),$
[…]
Discriminative Perceptual Transfer. With the discriminative category proxy set C, and the retrieval embeddings $\hat{\textbf{E}}^{R}$ of the origin training images are used for distillation:
(8) $\mathcal{L}_{dpt}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}\right).$
After optimizing $\mathcal{L}_{dpt}$, the image encoder adapts to capture subtle inter-category differences transferred from the discriminative branch. This enhancement enables effective retrieval of visually similar items without relying on the computationally expensive discriminative perception module.
3.4. Overall Function
As illustrated in Fig. 2, our objective function comprises two components: the proxy learning loss and the distillation loss $\mathcal{L}_{dpt}$. Finally, the total loss for our framework can be defined as:
(9) $\mathcal{L}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right)+\beta\mathcal{L}_{dpt},$
where $\textbf{C}\in\{\textbf{C},\textbf{C}^{\mathrm{b}}\}$ contains original training category and background category, and $\beta$ is the hyper-parameter to weight the $\mathcal{L}_{dpt}$.
4. Experiments
Table 2. Comparison with state-of-the-art methods in the closed-set setting on CUB-200-2011 and Stanford Cars. The best result is shown in bold, and the second-best result is underlined.
CUB-200-2011 Stanford Cars
Method Backbone Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) CNN 80.4 85.7 89.3 92.3 77.2 83.0 87.2 90.2
HIST (Lim et al., 2022) CNN 75.6 83.0 88.3 91.9 89.2 93.4 95.9 97.6
PNCA++  (Teh et al., 2020) CNN 80.4 85.7 89.3 92.3 86.4 92.3 96.0 97.8
Hyp-ViT (Ermolov et al., 2022) ViT 84.2 91.0 94.3 96.0 76.7 85.2 90.8 94.7
DVA (Ours) ViT 85.2 91.4 94.6 96.2 88.0 93.2 97.0 98.5
Table 3. Comparison with state-of-the-art methods in the open-set setting on CUB-200-2011 and Stanford Cars.
CUB-200-2011 Stanford Cars
Method Backbone Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
PNCA  (Movshovitz-Attias et al., 2017) CNN 49.2 61.9 67.9 72.4 73.2 82.4 86.4 88.7
DGCRL (Zheng et al., 2019) CNN 67.9 79.1 86.2 91.8 75.9 83.9 89.7 94.0
DAS (Liu et al., 2022) CNN 69.2 79.3 87.1 92.6 87.8 93.2 96.0 97.9
IBC (Seidenschwarz et al., 2021) CNN 70.3 80.3 87.6 92.7 88.1 93.3 96.2 98.2
Proxy-Anchor (Kim et al., 2020) CNN 71.1 80.4 87.4 92.5 88.3 93.1 95.7 97.0
HIST (Lim et al., 2022) CNN 71.4 81.1 88.1 - 89.6 93.9 96.4 -
PNCA++  (Teh et al., 2020) CNN 70.1 80.8 88.7 93.6 90.1 94.5 97.0 98.4
FRPT (Wang et al., 2023b) CNN 74.3 83.7 89.8 94.3 91.1 95.1 97.3 98.6
DFML (Wang et al., 2023c) ViT 79.1 86.8 - - 89.5 93.9 - -
Hyp-ViT (Ermolov et al., 2022) ViT 84.0 90.2 94.2 96.4 86.0 91.9 95.2 97.2
DVA (Ours) ViT 84.9 90.6 94.5 96.7 90.7 94.8 97.1 98.4
Table 4. Comparison with state-of-the-art methods on Stanford Dogs.
Open-set Setting Closed-set Setting
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 86.1 91.9 95.7 97.6 83.7 89.9 93.9 96.5
HIST (Lim et al., 2022) 86.2 92.3 95.6 97.5 84.7 90.1 93.4 95.6
PNCA++  (Teh et al., 2020) 86.0 92.3 95.6 97.7 83.8 90.1 94.7 96.8
Hyp-ViT (Ermolov et al., 2022) 87.8 92.8 95.9 97.6 79.2 86.9 91.8 95.2
DVA (Ours) 90.4 94.6 97.0 98.3 87.7 93.0 95.8 97.3
4.1. Experiment Setup
4.1.1. Datasets
CUB-200-2011 (Branson et al., 2014) comprises 11, 788 bird images from 200 bird species. In the closed-set setting, the dataset is divided into training and testing subsets comprising 5,994 and 5,794 images, respectively, out of a total of 11,788 images. For the open-set setting, we employ the first 100 subcategories (comprising 5,864 images) for training, and the remaining subcategories (comprising 5,924 images) are used for testing.
Stanford Cars (Krause et al., 2013) consists of 16,185 images depicting 196 car variants. Similarly, these images were split into 8,144 training images and 8,041 test images in the closed-set setting. For the open-set setting, we utilize the first 98 subcategories (comprising 8,054 images) for training and the remaining 98 subcategories (comprising 8,131 images) for testing.
Stanford Dogs (Dataset, 2011) contains 20,580 images showcasing dogs across 120 subcategories. We use the first 60 subcategories with 12,000 images for training and the remaining 60 subcategories with 8580 images for testing for the closed-set setting. For the open-set setting, we utilize the first 60 subcategories (comprising 10,651 images) for training and the remaining 60 subcategories (comprising 9,929 images) for testing.
NABirds (Horn et al., 2015) contains 48,562 images showcasing North American birds across 555 subcategories. We use the entire dataset as the generalization test dataset.
FGVC Aircraft (Maji et al., 2013) contains 10,000 images showcasing 100 subcategories. We use the entire dataset as the generalization test.
INat2021 (Van Horn et al., 2021) contains more than 2.7 million images of 10,000 categories. For the convenience of evaluation, we use the 27,105 images from the first 100 subcategories in its training set as the generalization test dataset.
4.1.2. Evaluation protocols
To evaluate retrieval performance, we adopt Recall@K with cosine distance in previous work (Song et al., 2016), which calculates the recall scores of all query images in the test set. For each query image, the top K similar images are returned. A recall score of 1 is assigned if at least one positive image among the top K images; otherwise, it is 0.
4.1.3. Implementation details
In experiments, we employ the ViT-B-16 (Dosovitskiy et al., 2021) pre-trained on ImageNet21K (Russakovsky et al., 2015) as our image encoder. All input images are resized to 256 $\times$ 256, and crop them into 224 $\times$ 224. In the training stage, we utilize the Adam optimizer with weight decay of 1e-4, and employ cosine annealing as the optimization scheduler. The learning rate for all datasets is initialized to 1e-1 except for Stanford Dogs, which has a learning rate of 1e-2. The number of training epochs for all datasets is set to 10 and the batch size is set to 32. We train our model on a single NVIDIA 3090 GPU to accelerate the training process.
4.2. Comparison with State-of-the-Art Methods
4.2.1. Closed-set Setting
We first compare our proposed DVA with previous competitive methods under closed-set setting. Quantitative comparison results for the CUB-200-2011 and Stanford Cars datasets are presented in Table 2, and the results for the Stanford Dogs dataset can be found in Table 4. From these tables, it can be observed that our proposed DVA outperforms other state-of-the-art methods on CUB-200-2011 and Stanford Dogs, and achieves competitive performance on Stanford Cars, which demonstrates the enhanced discriminative capability of our DVA for fine-grained visual retrieval. Specifically, in comparison with Hyp-ViT (Ermolov et al., 2022), the current state-of-the-art on CUB-200-2011, our DVA demonstrates a 1.0% improvement in Recall@1. The experimental results on the Stanford Cars indicate that our method outperforms the most of existing methods but falls slightly behind HIST (Lim et al., 2022). We argue the possible reason may be attributed to the relatively regular and simpler shape of the cars.
4.2.2. Open-set Setting
The open-set setting poses greater challenges compared to the closed-set setting due to the unknown subcategories in the test data. The experimental results for the CUB-200-2011 and Stanford Cars datasets are shown in Table 3, while the results for the Stanford Dogs dataset are provided in Table 4. The results reveal a consistent trend in both open and closed-set settings: our method outperforms other state-of-the-art approaches on CUB-200-2011 and Stanford Dogs, while delivering competitive results on the Stanford Cars dataset. To be specific, in comparison with Hyp-ViT (Ermolov et al., 2022) on CUB-200-2011, our DVA exhibits a 0.9% improvement in Recall@1. Experimental results on the Stanford Cars dataset show that our method outperforms most existing methods but lags behind FRPT (Wang et al., 2023b) by a slight margin. The results of experiments on Stanford Dogs are shown in Table 4. Our DVA demonstrates a clear performance advantage over Hyp-ViT, further indicating its superior generalizability.
4.3. Generalizability Analysis
To further demonstrate the generalization ability of our method, we conduct comparative experiments on the out-of-distribution datasets NABirds (Horn et al., 2015), FGVC Aircraft (Maji et al., 2013), and INat2021 (Van Horn et al., 2021). The results are presented in Tables 5, 6, and 7, respectively. Our method consistently outperforms all other approaches across these datasets. Notably, it demonstrates a clearer advantage compared to the results discussed in Section 4.2, highlighting its strong generalization capability. These findings confirm that DVA not only distinguishes similar samples within the distribution effectively, but also performs well on similar samples outside the distribution. This robust generalization ability is crucial for real-world applications where data distributions can vary significantly.
Table 5. Comparison on the NABirds dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 51.0 61.3 70.3 77.9 19.6 26.5 34.6 44.0
HIST (Lim et al., 2022) 50.4 60.3 69.3 76.8 14.4 20.1 26.4 34.4
PNCA++  (Teh et al., 2020) 57.9 68.7 77.5 84.7 36.5 46.5 56.9 66.9
Hyp-ViT (Ermolov et al., 2022) 64.7 74.6 82.7 88.7 44.6 55.3 65.3 74.3
DVA (Ours) 79.2 86.4 91.5 94.8 70.5 80.1 87.2 92.1
Table 6. Comparison on the FGVC Aircraft dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 28.7 37.6 46.9 57.0 32.3 41.4 51.3 62.0
HIST (Lim et al., 2022) 31.4 39.9 49.3 59.3 32.4 41.3 50.6 60.8
PNCA++  (Teh et al., 2020) 38.9 48.8 58.8 69.2 36.9 47.1 57.3 67.8
Hyp-ViT (Ermolov et al., 2022) 39.3 49.3 59.3 69.4 42.6 52.2 62.5 72.0
DVA (Ours) 42.2 52.1 62.2 72.3 48.3 56.2 68.1 77.2
Table 7. Comparison on the INat2021 dataset with state-of-the-art methods trained on the CUB-200-2011 and Stanford Cars datasets.
CUB-200-2011 Standford Cars
Method Recall@1 Recall@2 Recall@4 Recall@8 Recall@1 Recall@2 Recall@4 Recall@8
Proxy-Anchor (Kim et al., 2020) 24.0 33.4 44.0 55.4 13.7 20.9 30.1 41.5
HIST (Lim et al., 2022) 21.1 29.9 40.5 52.0 13.4 20.1 29.2 40.5
PNCA++  (Teh et al., 2020) 30.1 40.7 52.0 64.0 25.6 35.5 47.3 59.4
Hyp-ViT (Ermolov et al., 2022) 38.4 50.2 62.1 72.6 31.7 43.0 55.1 66.9
DVA (Ours) 51.8 63.7 73.9 82.3 45.0 57.2 68.3 78.2
4.4. Ablation Studies and Analysis
Table 8. The Recall@K results (%) of component ablation study on CUB-200-2011.
Setting R@1 R@2 R@4 Latency (ms)
Base 72.5 82.6 89.1 1.51
Base + ICA 83.6 90.3 93.9 1.51
Base + ICA + OPA 86.8 92.0 94.7 3.65
Base + ICA + OPA + DPT 84.9 90.6 94.5 1.51
4.4.1. Efficacy of various components
The proposed DVA comprises three essential components: Object-Perceptual Adaption (OPA), In-Context Adaptation (ICA), and Discriminative Perceptual Transfer (DPT). We conducted ablation experiments on these components, and the results are reported in Table 8, with the Base representing the pure ViT. The introduction of ICA improves the Recall@1 accuracy by 11.1% on CUB-200-2011, which shows that ICA effectively guides the FGIR task to the task solved in the original process through feature adjustment. OPA further improves the ability of the pre-trained model to distinguish similar subcategories on FGIR, bringing a gain of 3.2% Recall@1 on CUB-200-2011, but it brings huge computational overhead. DPT effectively removes the huge computational overhead brought by OPA and achieves a balance between performance and retrieval efficiency.
4.4.2. Sensitivity analysis of parameter $\beta$
Figure 3. Analyses of hyper-parameter $\beta$ on CUB-200-2011.
As illustrated in Fig. 3, we investigate the impact of varying values of $\beta$, representing the weight of loss in Eq. (9). The performance gradually increases with the value but decreases when the value exceeds 3. The drop in performance can be attributed to the fact that as the value of $\beta$ increases, the learned feature space tends to receive samples adjusted by the OPA module, while the input at inference is not processed by the OPA, resulting in a difference in the input space that causes performance degradation. Consequently, we set $\beta$ to 3 for all datasets.
Figure 4. Examples of top-10 retrieved images on CUB-200-2011 by DVA. The images with green boxes are the correct ones, and those with red boxes are the wrong ones.
Table 9. The Recall@K results (%) of component ablation study on CUB-200-2011.
Setting R@1 R@2 R@4 R@8 Prams (M)
$\textbf{IC}_{q}$ 82.7 89.1 92.9 95.6 0.295
$\textbf{IC}_{k}$ 81.8 88.8 93.1 95.5 0.295
$\textbf{IC}_{v}$ 84.0 90.0 93.6 95.6 0.295
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$ 84.9 90.6 94.5 96.7 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{v}$ 84.5 90.5 93.9 95.7 0.590
$\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 82.6 89.0 93.4 95.4 0.590
$\textbf{IC}_{q}$, $\textbf{IC}_{k}$, $\textbf{IC}_{v}$ 84.4 90.2 93.8 95.9 0.885
4.4.3. Analysis of In-Context Adaptation
To validate the effectiveness of the proposed ICA components, we further divide the ICA module into three parts: $\textbf{IC}_{q}$, $\textbf{IC}_{k}$, and $\textbf{IC}_{v}$, corresponding to the Q, K, and V projectors in the Attention Projector, respectively. We perform ablation studies on these components, and the results are presented in Table 9. The results show that using $\textbf{IC}_{q}$ alone or in combination with $\textbf{IC}_{k}$ and $\textbf{IC}_{v}$ leads to better performance. Notably, the best results are achieved when $\textbf{IC}_{q}$ and $\textbf{IC}_{k}$ are used together. Therefore, we adopt this combination to construct the final ICA module in our DVA.
4.4.4. Visualization
After visualizing the top 10 retrieval results on CUB-200-2011 in Fig. 4, we observe that DVA effectively retrieves images from the same subcategory, even across different variations and backgrounds. This demonstrates that our dual-vision adaptation mechanism enhances the pre-trained model’s ability to focus on fine-grained characteristics. For the failure case in the second row, even humans can easily overlook the subtle differences between the query image and the returned image and make mistakes. To explore the superiority of the designed DVA, We also conduct visualization as shown in Fig. 5. The first column displays the input images. The second column shows the class activation maps generated by GradCAM for the baseline model, which processes the input images. The third column presents the class activation maps for DVA using the same input. The results indicate that DVA effectively reduces attention to background regions and improves the baseline model’s focus on critical areas, capturing finer details in subcategories such as the head, wings, and tail.
Figure 5. Class activation visualizations on CUB-200-2011. For each sample, we show the input image, the class activation map of the baseline model, and the proposed DVA.
4.4.5. Comparison with other efficient parameter fine-tuning methods
We compare our method with two efficient parameter fine-tuning approaches: CLIP-Adapter (Gao et al., 2024) and AdaptFormer (Chen et al., 2022a). As shown in Table 10, both methods perform well on the CUB-200-2011 dataset. However, our DVA still achieves superior results, demonstrating its ability to capture subtle differences and accurately distinguish similar subcategories through dual-vision adaptation. Furthermore, DVA outperforms the fully fine-tuned ViT. This is expected, as fine-tuning a pre-trained model on a small fine-grained dataset often causes overfitting, reducing its generalization capability.
Table 10. The Recall@K results (%) of different parameter efficient fine-tuning methods on CUB-200-2011.
Methods R@1 R@2 R@4 R@8 Params (M)
Fine-tuning ViT 82.7 88.7 92.5 95.2 85.6
CLIP-Adapter (Gao et al., 2024) 69.8 80.0 87.5 92.1 0.295
AdaptFormer (Chen et al., 2022a) 73.3 83.4 89.3 93.2 0.590
DVA (Ours) 84.9 90.6 94.5 96.7 0.590
5. Conclusion
In this paper, we propose Dual-Vision Adaptation (DVA), a method aimed at improving generalization in fine-grained image retrieval. While conventional methods risk overfitting the training set due to full-model fine-tuning, DVA mitigates this issue by preserving pre-trained knowledge. DVA introduces Object-Perceptual Adaption (OPA) and In-Context Adaptation (ICA) to guide a frozen pre-trained vision model in performing fine-grained retrieval tasks. Specifically, OPA leverages a visual foundation model to create a new background category and enhance pixels contributing to category prediction, and helping the model better perceive critical objects and their distinguishing features. ICA introduce a small set of in-context adaptation parameters that enables the frozen pre-trained model to capture discriminative details through feature adaptation. Additionally, DVA achieves a balance between discriminative perception and retrieval efficiency with the design of Discriminative Perceptual Transfer (DPT). Extensive experiments on three in-distribution and three out-of-distribution fine-grained benchmarks show that our DVA achieves excellent performance with very few learnable parameters.